how to draw water and oil projection which there are 2 containers and have hole at the same level?


1
Expert's answer
2021-07-09T08:31:50-0400

Let the density of the water is "\rho" and density of the oil is "\sigma"

Let the height of the hole from the ground is h,

Then "v^2=u^2+2gh"

here initial velocity in the horizontal direction = 0

"v = \sqrt{2gh}"

Both will have the same velocity at just before hitting the ground.

Applying the bernaull's equation,

"\rho g (H-h)=\frac{1}{2}\rho v^2"

So, it is not depending on the density of liquid.